OpenStudy (anonymous):

In the function f(t) = sin(t), what is the effect of multiplying t by a coefficient of 2?

OpenStudy (anonymous):

The period is reduced by 2.

OpenStudy (anonymous):

a. changed amp to 2 b. changed amp to 1/2 c. changed period to pi d. changed period to 4pi

OpenStudy (anonymous):

The period of f(t)=sin(t) is 2*Pi, now the period of f(2t)=sin(2t) is reduced by 2, which is ....

OpenStudy (anonymous):

drawar is right and if you know hat the one period of \( \sin(x)\) is \(2\pi\) you can do the magic.

OpenStudy (anonymous):

I'm not so bright with this. I've got this answer wrong serveral times, hahah.

OpenStudy (anonymous):

So do I need to plug it into the calculator to figure out where it changes???

OpenStudy (anonymous):

If you're not good with visualizing functions, i can recommend geogebra. It's a rather easy to use function plotter which is free. You can get it at http://www.geogebra.org/cms/ Now back to your problem. The normal period is \(2\pi\) so if it is halved by \(2\) you can simply calculate it by \(\frac{2\pi}{2} = \pi\) as the two in the denominator and enumerator cancels.

OpenStudy (anonymous):

So basically it's asking what will happen when you add pi to it?

OpenStudy (anonymous):

So if i get your problem right you got \(f(x) = \sin(x)\) as your function you have to explain what changes if you have a look \(g(x) =\sin(2x)\), is that right?

OpenStudy (anonymous):

O: I think so

OpenStudy (anonymous):

Hang on i'll do a quick plot to make things clearer.

OpenStudy (anonymous):

NO WAIT IT'S D

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

... Right?

OpenStudy (anonymous):

No, actually D is wrong

OpenStudy (anonymous):

What...

OpenStudy (anonymous):

No because it goes from 9pi/4 to 13pi/4

OpenStudy (anonymous):

So the blue curve is \(\sin(x)\) and the black one \(\sin(2x)\)

OpenStudy (anonymous):

You can see that both amplitudes are equally high so there's no change in that.

OpenStudy (anonymous):

If we have a look at the period we can see that it takes \(2\pi\) for \(\sin(x)\) to start over. Whereas \(\sin(2x)\) starts repeating after it reaches \(\pi\)

OpenStudy (anonymous):

I saw...

OpenStudy (anonymous):

Is it clear so far or do you still have trouble understanding it?

OpenStudy (anonymous):

i dont understandhow im wrong.

OpenStudy (anonymous):

Tell me how you got answer D

OpenStudy (anonymous):

I did.